# Flow-GRPO: Training Flow Matching Models via Online RL

### v2

@@ 5 Experiments
In this section, we empirically evaluate Flow-GRPO’s ability to improving flow matching models on three representative tasks. (1) Composition Image Generation: This task requires precise object arrangement and attribute control. We report the results on GenEval. (2) Visual Text Rendering: a rule-based task that stresses the accurate rendering of the text specified in the prompt. (3) Human Preference Alignment: This task aims to align T2I models with human preferences. We use PickScore [19] as our reward model.
@@ 5.1 Experimental Setup
We introduce three tasks, detailing their respective prompts and reward definitions. For hyperparameter details and compute resource specifications, please refer to Appendix B.2 and Appendix B.2.1.
Compositional Image Generation
@@
  • Counting: $r=|N_{\text{gen}}-N_{\text{ref}}|/N_{\text{ref}}$.
  • Position / Color: If the object count is correct, a partial reward is assigned; the remainder is granted when the predicted position or color is also correct.
@@
Visual Text Rendering [8]
Text is common in images such as posters, book covers, and memes, so the ability to place accurate and coherent text inside the generated images is crucial for T2V models. In our settings, we define an OCR task, where each prompt follows the template ‘‘A sign that says ‘‘text’’. Specifically, the placeholder ‘‘text’’ is the exact string that should appear in the image. We use GPT to produce 20K training prompts and 1K test prompts. Following [65], we measure text fidelity with the reward
$r=\frac{|N_{\text{gen}}-N_{\text{ref}}|}{N_{\text{ref}}},$
where $N_{\text{gen}}$ is the number of characters detected in the image and $N_{\text{ref}}$ is the number of characters inside the quotation marks in the prompt. This reward also serves as our metric of text accuracy.
Human Preference Alignment [19]
This task aims to align T2I models with human preferences. We use PickScore  [^2] https://huggingface.co/yuvalkirstain/PickScore_v1 as our reward model, which is based on large-scale human annotated pairwise comparisons of images generated from the same prompt. For each image and prompt pair, PickScore provides an overall score that evaluates multiple criteria, such as the alignment of the image with the prompt and its visual quality.
Image Quality Evaluation Metric
Because the T2I model is trained to maximise a predefined reward signal in RL, it is susceptible to reward hacking—situations in which the numerical reward rises but the quality or diversity of the generated images declines. This study aims to make online RL truly effective for T2I generation without compromising quality or diversity. To detect potential reward hacking beyond task-specific accuracy, we evaluate four automatic image quality metrics:
All metrics are evaluated on DrawBench [1], a comprehensive and challenging benchmark with diverse prompts for T2I models.
@@ 5.2 Main Results
Figure 1 and Table 1 show Flow-GRPO’s GenEval performance steadily improving during training, ultimately outperforming GPT-4o. This occurs while maintaining both image quality metrics and preference scores on DrawBench, a benchmark with diverse and comprehensive prompts for evaluating general model capabilities. Figure 3 offers qualitative comparisons. Beyond Compositional Image Generation, Table 2 details evaluations on Visual Text Rendering and Human Preference tasks. Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench. See Figures 8 and 9 in Appendix C.3 for related qualitative examples. For the Human Preference task, image quality did not decrease without KL regularization. However, we found that omitting KL caused a collapse in visual diversity—a form of reward hacking discussed further in Section 5.3. These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
@@ 5.3 Analysis
Reward Hacking
We explored two approaches to mitigate reward hacking: (1) Multi-reward ensemble: We combined preference-based rewards (e.g., UnifiedReward, PickScore, ImageReward), image quality rewards (e.g., Aesthetic Score, DeQA) with task reward. However, this also led to local blurriness and reduced diversity. Due to the challenge of covering all aspects of generation with diverse reward models, we abandoned this approach. (2) KL constraint: We tuned the KL coefficient to maintain a small, constant KL divergence during training, keeping the model close to its pretrained weight. This allows for task-specific reward optimization without harming general performance. As shown in Table 2, removing the KL constraint boosts performance on Compositional Image Generation and Visual Text Rendering but causes a sharp decline in image quality and preference scores on DrawBench. In contrast, a properly tuned KL constraint preserves quality while maintaining similar gains on task-specific metrics. In the Human Preference Alignment task, dropping the KL constraint doesn’t hurt image quality—likely due to PickScore overlap with evaluation metrics—but causes a collapse in visual diversity. Outputs converge to a single style, with different seeds producing nearly identical results. KL regularization prevents this collapse and preserves diversity. See Figure 7 in the Appendix C.2 for learning curves with and without KL, and Figure 4 for examples.
Figure 4: Effect of KL Regularization. The KL penalty effectively suppresses reward hacking, preventing potential Quality Degradation (for GenEval and OCR) and Diversity Decline (for PickScore).
Effect of Denoising Reduction
Figure 5 (a) highlights Denoising Reduction’s significant impact on accelerating training. To explore how different timesteps affect optimization, these experiments were conducted without the KL constraint. Reducing data collection timesteps from $40$ to $10$ achieved over a 4×4$\times$4 × speedup across all three tasks, without impacting final reward. While further tests showed this reduction didn’t consistently improve speed (and occasionally slowed some tasks), we standardized on $10$ timesteps for subsequent experiments, prioritizing the maintained reward. For the other two tasks, learning curves of reward versus training time are presented in Figure 6 in the Appendix C.1.
Effect of Noise Level
Higher $\sigma_{t}$ in the SDE boosts image diversity and exploration, vital for RL training. We control this exploration with a noise level $a$ (Eq. 12). Ablation studies (Figure 5 (b)) show $a$’s impact on performance. Small $a$ (e.g., 0.1) limits exploration and slows reward improvement. Increasing $a$ (up to 0.7) enhances exploration and accelerates reward gains. Beyond this, further increases (e.g., $a=0.7\rightarrow 1.0$) yield no extra benefit, as exploration is sufficient. However, too much noise degrades image quality, leading to zero reward and failed training. We advise using the highest noise level that maintains good image quality.
@@
Generalization Analysis
Flow-GRPO demonstrates strong generalization on unseen scenarios from GenEval (Table 4). Specifically, it captures object number, color, and spatial relations, generalizing well to unseen object classes. It also effectively controls object count, generalizing from training on $2-4$ objects to generate $5-6$ objects. Furthermore, Table 3 shows Flow-GRPO achieves significant gains on T2I-CompBench++. This comprehensive benchmark for open-world compositional T2I generation features object classes and relationships substantially different from our model’s GenEval-style training data.
@@
While this work focuses on T2I tasks, Flow-GRPO’s potential applicability to video generation [25, 27] raises several questions for future research: (1) Reward Design for Video: Defining effective reward models for video generation is crucial. Simple heuristics like object detectors or trackers can promote physical realism and temporal consistency, but more sophisticated models are needed. (2) Balancing Multiple Rewards: Effective video generation requires optimizing multiple objectives (e.g., realism, smoothness, coherence). Balancing these often conflicting signals is challenging and demands careful tuning. (3) Scalability: Video generation is significantly more resource-intensive than T2I. Scaling Flow-GRPO for video thus demands more efficient data collection and training pipelines.
@@ Appendix A Mathematical Derivations for Stochastic Sampling using Flow Models
We consider its stochastic counterpart. Inspired by the derivation from SDE to its probability flow ODE in SGMs [23], we aim to construct a SDE with specific drift and diffusion coefficients so that its marginal distribution matches that of Eq. 14. We begin with the generic form of SDE:
@@
$-\nabla\cdot[f_{\text{SDE}}p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}p_{t}({\bm{x}})]$ (18)
@@
Substituting Eq. 19 to Eq. 18, we arrive at the drift coefficients of the target SDE:
$f_{\text{SDE}}={\bm{v}}_{t}-\frac{1}{2}\sigma_{t}^{2}\nabla\log p_{t}({\bm{x}})$ (20)
Hence we can rewrite the reverse-time SDE in Eq. 15 as:
@@
where $d{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ is the diffusion coefficient that controls the level of stachasticity during sampling.
Once the score $\nabla\log p_{t}({\bm{x}}_{t})$ is available, the stochastic process can be sampled directly. While the Flow-Matching framework, this score is tied to the velocity field ${\bm{v}}_{t}$.
@@ Appendix B Further Details on the Experimental Setup
B.1 Model Specification
@@ B.2 Hyperparameters Specification
B.2.1 Compute Resources Specification
@@ Appendix C Extended Experimental Results
C.1 Effect of Denoising Reduction
@@
C.2 Learning Curves with or without KL
@@
C.3 Qualitative Results
In addition to the GenEval reward comparisons in Figure 3, Figures 8 and 9 show qualitative comparisons between SD-3.5-M and SD-3.5-M + Flow-GRPO trained with PickScore and OCR rewards, respectively.
Figure 8: Qualitative comparison between the SD-3.5-M and SD-3.5-M + Flow-GRPO trained with PickScore reward.
Figure 9: Qualitative comparison between the SD-3.5-M and SD-3.5-M + Flow-GRPO trained with OCR reward.

### v3

@@ 5 Experiments
This section empirically evaluates Flow-GRPO’s ability to improve flow matching models on three tasks. (1) Composition Image Generation: This task requires precise object arrangement and attribute control. We report the results on GenEval. (2) Visual Text Rendering: a rule-based task that evaluates the accurate rendering of the text specified in the prompt. (3) Human Preference Alignment: This task aims to align T2I models with human preferences.
@@ 5.1 Experimental Setup
We introduce three tasks, detailing their respective prompts and reward definitions. For hyperparameter details and compute resource specifications, please refer to Appendix B.3 and Appendix B.4.
Compositional Image Generation.
@@
Visual Text Rendering [8].
Text is common in images such as posters, book covers, and memes, so the ability to place accurate and coherent text inside the generated images is crucial for T2I models. In our settings, we define an text rendering task, where each prompt follows the template ‘‘A sign that says ‘‘text’’. Specifically, the placeholder ‘‘text’’ is the exact string that should appear in the image. We use GPT4o to produce 20K training prompts and 1K test prompts. Following [56], we measure text fidelity with the reward $r=\max(1-N_{\text{e}}/N_{\text{ref}},0),$ where $N_{\text{e}}$ is the minimum edit distance between the rendered text and the target text and $N_{\text{ref}}$ is the number of characters inside the quotation marks in the prompt. This reward also serves as our metric of text accuracy.
Human Preference Alignment [19].
This task aims to align T2I models with human preferences. We use PickScore [19] as our reward model, which is based on large-scale human annotated pairwise comparisons of images generated from the same prompt. For each image and prompt pair, PickScore provides an overall score that evaluates multiple criteria, such as the alignment of the image with the prompt and its visual quality.
Image Quality Evaluation Metric.
Since the T2I model is trained to maximize a predefined reward, it is vulnerable to reward hacking, where the reward increases but image quality or diversity declines. This study aims to make online RL effective for T2I generation without noticeably compromising quality or diversity. To detect reward hacking beyond task-specific accuracy, we evaluate four automatic image quality metrics: Aesthetic Score [57], DeQA [58], ImageReward [32], and UnifiedReward [59] (see Appendix B.1 for details). All metrics are computed on DrawBench [1], a comprehensive benchmark with diverse prompts for T2I models.
@@ 5.2 Main Results
Figure 1 and Table 1 show Flow-GRPO’s GenEval performance steadily improving during training, ultimately outperforming GPT-4o. This occurs while maintaining both image quality metrics and preference scores on DrawBench, a benchmark with diverse and comprehensive prompts for evaluating general model capabilities. Figure 3 offers qualitative comparisons. Beyond Compositional Image Generation, Table 2 details evaluations on Visual Text Rendering and Human Preference tasks. Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench. See Figures 9, 10 & 11 in Appendix C.4 for related qualitative examples. For the Human Preference task, image quality did not decrease without KL regularization. However, we found that omitting KL caused a collapse in visual diversity, a form of reward hacking discussed further in Section 5.3. These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
We also compare Flow-GRPO with other alignment methods, including supervised fine-tuning, reward weighted regression [14, 69], Flow-DPO [14], and their online variants. Flow-GRPO outperforms all of them by a large margin. See Appendix C.1 for details.
@@ 5.3 Analysis
Reward Hacking.
We use KL regularization to mitigate reward hacking by tuning the KL coefficient to keep the divergence small and nearly constant during training, keeping the model close to its pretrained weights. This allows task-specific reward optimization without harming overall performance. As shown in Table 2, removing the KL constraint for Compositional Image Generation and Visual Text Rendering significantly reduces image quality and preference scores on DrawBench. In contrast, a properly tuned KL preserves quality while achieving similar gains on task-specific metrics. In the Human Preference Alignment task, removing KL does not affect image quality, likely due to overlap between PickScore and evaluation metrics, but causes a collapse in visual diversity. Outputs converge to a single style, with different seeds producing nearly identical results. KL regularization prevents this collapse and maintains diversity. See Figure 8 in Appendix C.3 for training curves and Figure 4 for more examples.
Figure 4: Effect of KL Regularization. The KL penalty effectively suppresses reward hacking, preventing Quality Degradation (for GenEval and OCR) and Diversity Decline (for PickScore).
Effect of Denoising Reduction.
Figure 5 (a) highlights Denoising Reduction’s significant impact on accelerating training. To explore how different timesteps affect optimization, these experiments are conducted without the KL constraint. Reducing data collection timesteps from $40$ to $10$ achieves over a 4×4$\times$4 × speedup across all three tasks, without impacting final reward. Further reducing to $5$ does not consistently improve speed and sometimes slows training, so we choose $10$ timesteps for later experiments. For the other two tasks, learning curves of reward versus training time are presented in Figure 7 in the Appendix C.2.
Effect of Noise Level.
Higher $\sigma_{t}$ in the SDE boosts image diversity and exploration, vital for RL training. We control this exploration with a noise level $a$ (Eq. 9). Figure 5 (b) shows the impact of $a$ on performance. A small $a$ (e.g., 0.1) limits exploration and slows reward improvement. Increasing $a$ (up to 0.7) boosts exploration and speeds up reward gains. Beyond this point (e.g., from 0.7 to 1.0), further increases provide no additional benefit, as exploration is already sufficient. We also observe that injecting too much noise by further increasing $a$ degrades image quality, resulting in zero reward and failed training.
@@
Generalization Analysis.
Flow-GRPO demonstrates strong generalization on unseen scenarios from GenEval (Table 4). Specifically, it captures object number, color, and spatial relations, generalizing well to unseen object classes. It also effectively controls object count, generalizing from training on $2-4$ objects to generate $5-6$ objects. Furthermore, Table 3 shows Flow-GRPO achieves significant gains on T2I-CompBench++ [6, 70]. This comprehensive benchmark for open-world compositional T2I generation features object classes and relationships substantially different from our model’s GenEval-style training data.
@@
6 Conclusion
We have presented Flow-GRPO, the first method to integrate online RL into flow matching models. By converting deterministic ODEs to SDEs and reducing denoising steps during training, Flow-GRPO enables efficient RL-based optimization without noticeably compromising image quality or diversity. Our method significantly improves performance on compositional generation, text rendering, and human preference alignment, with minimal reward hacking. Flow-GRPO offers a simple and general framework for applying online RL to flow-based generative models.
@@ 6 Conclusion
Although this work focuses on T2I tasks, Flow-GRPO has potential for video generation [25, 27], raising several future directions: (1) Reward Design: Simple heuristics, such as using object detectors or trackers as rule-based rewards, can encourage physical realism and temporal consistency, but more advanced reward models are needed. (2) Balancing Multiple Rewards: Video generation requires optimizing multiple objectives, including realism, smoothness, and coherence. Balancing these competing goals remains challenging and demands careful tuning. (3) Scalability: Video generation is far more resource-intensive than T2I, so applying Flow-GRPO at scale requires more efficient data collection and training pipelines. Additionally, better methods for preventing reward hacking are worth exploring. While KL regularization helps significantly, it requires longer training and occasional reward hacking occurs for certain prompts.
@@ Appendix A Mathematical Derivations for Stochastic Sampling using Flow Models
We consider its stochastic counterpart. Inspired by the derivation from SDE to its probability flow ODE in SGMs [23], we aim to construct a forward SDE with specific drift and diffusion coefficients so that its marginal distribution matches that of Eq. 10. We begin with the generic form of SDE:
@@
$-\nabla\cdot[f_{\text{SDE}}\,p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}({\bm{x}}_{t},t)p_{t}({\bm{x}})]$ (14)
@@
Substituting Eq. 15 to Eq. 14, we arrive at the drift coefficients of the target forward SDE:
$f_{\text{SDE}}={\bm{v}}_{t}({\bm{x}}_{t},t)+\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}})$ (16)
Hence, we can rewrite the forward SDE in Eq. 11 as:
$\mathrm{d}{\bm{x}}_{t}=\left({\bm{v}}_{t}({\bm{x}}_{t})+\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})\right)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (17)
where $\mathrm{d}{\bm{w}}$ denotes Wiener process increments, and $\sigma_{t}$ is the diffusion coefficient controlling the level of stochasticity during sampling.
The relationship between forward and reverse-time SDEs has been established in [72, 23]. Specifically, if the forward SDE takes the form
$\mathrm{d}{\bm{x}}_{t}=f({\bm{x}}_{t},t)\,\mathrm{d}t+g(t)\,\mathrm{d}{\bm{w}},$ (18)
then the corresponding reverse-time SDE is
$\mathrm{d}{\bm{x}}_{t}=\left[f({\bm{x}}_{t},t)-g^{2}(t)\nabla\log p_{t}({\bm{x}}_{t})\right]\,\mathrm{d}t+g(t)\,\mathrm{d}\overline{{\bm{w}}}.$ (19)
Setting $g(t)=\sigma_{t}$, we obtain the reverse-time SDE corresponding to Eq. 17 as
$\mathrm{d}{\bm{x}}_{t}$ $=\left[{\bm{v}}_{t}({\bm{x}}_{t})+\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})-\sigma_{t}^{2}\nabla\log p_{t}({\bm{x}}_{t})\right]\mathrm{d}t+\sigma_{t}\mathrm{d}\overline{{\bm{w}}}.$ (20)
We thus arrive at the final form of the reverse-time SDE:
@@
Once the score function $\nabla\log p_{t}({\bm{x}}_{t})$ is available, the process can be simulated directly. For flow matching, this score is implicitly linked to the velocity field ${\bm{v}}_{t}$.
@@ Appendix B Further Details on the Experimental Setup
B.1 Quality Metrics
The details of quality metrics are as follows:
B.2 Model Specification
@@ B.3 Hyperparameters Specification
B.4 Compute Resources Specification
@@ Appendix C Extended Experimental Results
C.1 Flow-GRPO vs. Other Alignment Methods
We compare Flow-GRPO with several alignment methods: supervised fine-tuning (SFT), reward-weighted regression (Flow-RWR [14, 69]), Flow-DPO [14], and their online variants. Flow-GRPO consistently outperforms all baselines by a significant margin. At each step, we generate a group of images using the same group size as in Flow-GRPO. The only difference lies in the update rule:
  • SFT: Select the highest-reward image in each group and fine-tune on it.
  • Flow-RWR[14, 69]: Apply a softmax over rewards in each group and perform reward-weighted likelihood maximization.
  • Flow-DPO[14]: Use the highest-reward image in each group as the chosen sample and the lowest as the rejected, then apply the DPO loss.
Offline variants use a fixed pretrained model for data collection, while online variants update their data collection model every 40 steps. As shown in Figure 6, Flow-GRPO outperforms all other methods. The figure also indicates that DPO and SFT improve over time. In contrast, RWR does not, which aligns with experimental findings on RWR in [12]. Additionally, Online DPO surpasses offline DPO, aligning with [15]’s finding that online DPO performs better. For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse.
Figure 6: Comparison of Flow-GRPO and Other Alignment Methods.
C.2 Effect of Denoising Reduction
@@
C.3 Learning Curves with or without KL
@@
C.4 Additional Qualitative Results
Figures 9, 10 & 11 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively. Flow-GRPO with KL regularization improves the target capability while maintaining image quality and minimizing reward-hacking. Conversely, removing the KL constraint significantly degrades image quality and diversity.
Figure 9: Additional Qualitative comparison between the SD3.5-M and SD3.5-M + Flow-GRPO trained with GenEval reward.
Figure 10: Additional Qualitative comparison between the SD3.5-M and SD3.5-M + Flow-GRPO trained with OCR reward.
Figure 11: Additional Qualitative comparison between the SD3.5-M and SD3.5-M + Flow-GRPO trained with PickScore reward.
C.5 Evolution of Evaluation Images During Flow-GRPO Training
To better understand the training dynamics of our proposed Flow-GRPO framework, we visualize the evolution of generated samples corresponding to fixed evaluation prompts at regular intervals during training in Figure 121314. For consistency, all visualizations are produced using a 40-step ODE-based sampling schedule. These qualitative results provide a visual representation of how the model progressively improves its generation quality and alignment with task objectives over time.
Figure 12: We visualize the generated samples across successive training iterations during the optimization of SD3.5-Medium on the GenEval task.
Figure 13: We visualize the generated samples across successive training iterations during the optimization of SD3.5-Medium on the OCR task.
Figure 14: We visualize the generated samples across successive training iterations during the optimization of SD3.5-Medium on the PickScore task.
Appendix D Training Sample Visualization with Denoising Reduction
In this section, we compare images obtained with SDE sampling at various steps against those produced by ODE sampling, and offer an intuitive view of the denoising reduction strategy. Figure 15 presents SD3.5-Medium samples under four inference settings: (a) ODE sampling with 40 steps; (b) SDE sampling with 40 steps; (c) SDE sampling with 10 steps; (d) SDE sampling with 5 steps.
The 40-step ODE and SDE runs yield visually indistinguishable images, confirming that our SDE sampler preserves quality. Shortening the SDE schedule to 10 and 5 steps introduces conspicuous artifacts, like color drift and fine details blur. Contrary to expectation that such low-quality samples might hinder optimization. it actually do just the opposite and accelerate optimization. Because Flow-GRPO relies on relative preferences, it still extracts a useful reward signal, while the shorter trajectories signifactly cut wall-clock time. Consequently, Flow-GRPO with denoising reduction strategy converges more quickly on both layout-oriented benchmarks such as GenEval and quality-focused metrics such as PickScore, without sacrificing final performance.
Figure 15: Visualization of training samples under difference inference settings.